OpenStudy (anonymous):

Jesse received two offers to join a music program. If he is enrolled in the program for 3 months, which program should he choose? First Program: 5 free songs for joining and 1 free song every month Second Program: 1 free song for joining and 2 free songs per month The first program The second program Either − He will get the same number of free songs from both. Not enough information

OpenStudy (anonymous):

1st program = 5 +1(3) 2nd program= 1+2(3)

OpenStudy (anonymous):

alright so 5+1=6 and 1+2=3 and 3 times 2 is 6

OpenStudy (anonymous):

what? PEMDAS or BODMAS you do the multiplication first, then you add 5+1(3) 5+3 8

OpenStudy (anonymous):

15